# Survey in Characterization of Semantic Change

### v4

@@ 2.3 Review Criteria
Table 1: Search terms used to discover papers related to change characterization.
@@
“novel sense change”,“innovative meaning”,
“new sense identification”,“neologism identification”
@@ 3 Corpora for Semantic Change Analysis
Corpora for semantic change can be selected based on differences in time (diachronic studies), domain (academic fields (Schlechtweg2019AWO)), culture (social media (Carpuat2013SenseSpottingNL; Shoemark2019RoomTG), countries), or dialect (Takamura2017AnalyzingSC; Uban2021CrosslingualLO), provided the corpora share a sufficient vocabulary for comparison.
In diachronic studies, some methods divide a corpus into discrete time windows, striking a balance between the size of the window and the quantity of data it contains. In contrast, others use a continuous representation of time, providing the timestamp of each text directly to the model (Hu2019DiachronicSM; Rosenfeld2018DeepNM; Wang2021Word2FunMW). These corpora were designed for specific purposes. For example, corpora with discrete time bins enable a comparative perspective on meaning, while continuous time representations allow for an evolutionary view of the changes.
With respect to characterization, the collected papers show an increasing interest in diachronic corpora, with many proposing handcrafted rules for automatically annotating change categories (Tang2013SemanticCC). Few works rely on human annotation (Schlechtweg2017GermanIF). While these approaches have advantages and drawbacks, the data required to train large representation models is often insufficient from a computational perspective.
Most of the works reviewed here follow a specific formula for diachronic studies, one historical corpus ($C_{1}$) and a more modern corpus ($C_{2}$). Where splitting Corpus of Historical American English (COHA) Davies2012ExpandingHI, a diversified corpus with a span from 1810s to the 2000s, in two corpus is the standard choice. Recent state-of-the-art NLP models require training on billions or even trillions of tokens (He2021DeBERTaV3ID), but diachronic corpora are often on the scale of millions of tokens (Schlechtweg2020Semeval). Given that increasing the amount of historical training data is often infeasible, a viable approach is to develop better algorithms for learning from limited data. In Table 2 we present the table of diachronic corpus in the reviewed works.
@@
Corpus Name Description Statistics (Date, Size) CLMET A collection of diachronic English corpora covering various text types like literary works, letters treatise and others. 34 million words, spanning from 1710-1920, divided into three 70-year sub-periods. COHA A balanced diachronic corpus of American English compiled from diverse document genres. Over 400 million words from about 107,000 documents published from the 1810s to the 2000s. PDN A Chinese newspaper corpus used for its faithful recording of contemporary language and events. Spans from 1946 to 2004, with a yearly average of 10,886,017 tokens. GoogleBook The largest available historical corpus, compiled from about 4% of all ever-published books. Over 1TB of text data, containing about 0.3 trillion words and spanning from 1600 to 2009. Kubhist2 A sample of a Swedish historical newspaper archive. From 1790-1830 containing about 71 million tokens. DTA A high-quality collection of historical German texts selected for representativeness and genre balance. 1,022 texts from the period 1741-1900. LC A collection of British English texts, encompassing various domains such as religion, politics, and law. Approximately one million words from 1640-1740. BNC A collection of primarily written British English sources. One hundred million words, from the late 20th century. NYT A collection of newspaper articles from The New York Times archives. 13 million articles (1851-present). Helsinki Comprises texts spanning Old English, Middle English, and Early Modern English. The analysis focused on the Middle and Modern English. Approximately 1.1 million words. Old (1̃150), Middle (1150-1500), and Modern (1500-1710) Gutenberg The bulk of English language literary works available through the project’s website. 4034 separate documents consisting of over 290 million words.
Even with progress in low-resource learning fields like rapid adaptation (Neubig2018RapidAO), few-shot learning (Liu2022FewShotPF), or in-context learning (Min2022RethinkingTR), there is still a significant lack of freely available corpora for semantic change studies (Kutuzov2018DiachronicWE).
@@ 4.1.1 Dimension
To detect changes along the dimension axis ($\mathcal{D}$), Sagi2009SemanticDA proposed counting co-occurrences of words and phonemes and measuring the cluster density of the cosine distance on a decomposition of the term-frequency matrix. They assumed that Latent Semantic Analysis is appropriate for examining vectors representing the context of each term’s occurrence and estimated cohesiveness by measuring the density of these context vectors. While they explored the methodology for taxonomic change analysis, they state that the method in fact measures word polysemy.
Tang2013SemanticCC proposed a framework for semantic change computation based on the assumption that it is an inherently successive and diachronic process. They distinguished between the word level (overall meaning) and the sense level (individual senses), building their approach around these two notions. A key step involves first identifying a word’s senses in a corpus and then analyzing their diachronic distribution.
Emms2016DynamicGM proposed a generative conditional model to detect if a word has acquired a new sense (a semantic neologism). They modeled the word’s sense distribution using Gibbs sampling, applying the model to time-stamped text to infer how the probabilities of a word’s senses change over time. The authors obtain all usages for a word in a specific year ($C_{1}$) from Google N-gram Corpus and compare with the usage frequency for the following next year($C_{2}$), obtaining n-grams that change above a threshold.
Bochkarev2022NeuralNA used a neural network to predict whether a word is used as a named entity. With this approach, they constructed a temporal view of word usage by tracking occurrences as nouns versus non-nouns over time, aiming to detect if the word acquired a new meaning.
@@ 4.1.2 Relation
The work of Tang2013SemanticCC transforms the problem of relation identification into detecting change patterns in time series data. They analyze semantic change at the word and individual-sense levels, using rules to infer whether a word is used metonymically or metaphorically. A different approach computes the entropy of target words to identify if a word was used in a non-standard context (Schlechtweg2017GermanIF). The authors split the usage of words from the Deutsches Textarchiv (DTA) in two time periods (varying by word) and measure the increase in entropy as a proxy for metaphorization. According to the authors, a word’s higher entropy makes it more likely to be used in a newly created metaphorical sense. However, the authors do not differentiate this from the emergence of a new, unrelated sense. They use a corpus of German documents to validate their approach.
@@ 4.1.3 Orientation
Frequency metrics that detect co-occurrences with sets of “good” or “bad” sentiment words have been widely used to infer the amelioration and pejoration of a word over time, especially for neologisms (Cook2010AutomaticallyIC; Jatowt2014AFF). Similarly, Buechel2016FeelingsFT used seed words in German and English to track emotions captured in corpora. They represent these emotions in the Valence-Arousal-Dominance framework rather than simple positive/negative connotation. A limitation of the latter study is that it doesn’t track the evolution of orientation in words.
@@ 4.2.1 Dimension
These approaches assume that changes in the number of topics assigned to a word can indicate changes in dimension. Lau2012WordSI introduced the idea of novelty as the proportion of a word’s usages corresponding to senses in a focus corpus versus a reference corpus. The goal is to use word sense induction based on topic modeling to detect if a word is used with a new meaning.
Cook2013ALA extended this work by quantifying a topic’s usage, estimating that a previously ”culturally silent” topic that becomes frequent in a focus corpus is highly likely to contain neologisms. They introduced the concept of topic relevance to compute a ranked list of potential neologisms.
Figure 4: Figure adapted from Inoue2022InfiniteSA. The stacked bar plots represent the topics obtained over time for the words ‘coach’, ‘record’ and ‘power’ respectively. We can observe new senses emerging and becoming dominant.
Lau2014LearningWS improved the definition of novelty by adding the word’s frequency in the corpus, then proposed combining novelty and relevance to rank neologisms. Frermann2016ABM introduced SCAN, a Bayesian topic model where topics are preserved across adjacent time periods. Inoue2022InfiniteSA introduced InfiniteSCAN, an improvement over SCAN. While most topic models rely on a predefined number of topics, InfiniteSCAN uses a logistic stick-breaking process to adapt the number of topics dynamically. This modification allows it to track a varying number of senses, enabling to count each meaning to measure broadening and narrowing, as shown in Figure 4. The authors split Clean COHA in intervals of 20 years to represent the senses.
@@ 4.2.2 Relation
To the best of our knowledge, no methodology for tackling conceptual change with topic-based representation has been published, although a work uses it in a synchronic detection setting (Heintz2013AutomaticEO).
@@ 4.2.3 Orientation
To the best of our knowledge, no methodology approaching orientation change with topic-based representation has been published. Wankhade2022ASO use topic models in their workflow to refine data for a sentiment classifier, but topics are not directly used to detect change in feelings.
@@ 4.3.1 Dimension
Comparing the size and structure of a graph over time can provide rich information about a word’s semantic change. biemann2006chinese was among the first to use this approach to detect sense broadening. He created a graph of associated nouns from a corpus and then applied the Chinese whispers algorithm to find clusters. By comparing these clusters over time, he could infer whether a word had acquired a new meaning.
Mitra2015AnAA analyzed tweets, counting word co-occurrences in bi-grams to build a graph. They ran a graph-clustering algorithm (Chinese Whispers) to determine a set of senses, where each cluster represents a different sense. By comparing these sets of senses over time, they could infer if a word’s senses split or became broader.
Figure 5: Adaptation from Ehmller2020SenseTD. Ego-network, built from word co-occurrence graph, for ‘mouse’. We observe that in 1830 it was used with the sense of ‘weak’ and ‘rat,’ where in 1960 the sense of ‘computer device’ emerged.
Mathew2017Adapting improved the filtering and sense selection methods of previous works (Mitra2015AnAA; Mccarthy2004Finding; Lau2014LearningWS) to better capture sense evolution in Google Books and newspaper datasets. Tahmasebi2017FindingIW considered a graph of two-word noun phrases as a cluster of meaning, then merged sub-graphs based on Lin similarity. They tracked change in dimension by analyzing how these sub-graphs change over time.
Ehmller2020SenseTD built a co-occurrence graph, filtering it with point-wise mutual information to remove non-relevant edges, and then created ego-networks to run graph-clustering algorithms. This approach allowed them to build a sense tree—a tree of word usage evolution where new senses are new branches, as shown in Figure 5. Similar to previous work, the authors restricted their analysis to nouns only.
@@ 4.3.2 Relation
Graphs are a good way to represent word relations, and some approaches have used them to detect metonymy usage, e.g., Teraoka2016MetonymyAU; Pedinotti2020Dont. Their models use WordNet or a graph-based distributional model to determine a contextualized representation of word semantics. However, to the best of our knowledge, no work has proposed using graphs at a representation level to compare which words have changed their relations between corpora.
@@ 4.3.3 Orientation
The orientation of words is generally inferred from similarity or proximity to seed words. Many word graphs are constructed for this purpose, and some works (Reagan2017SentimentAM; Barnes2021StructuredSA) use these graphs to infer orientation. Hamilton2016InducingDS used the SentiProp algorithm to propagate a word’s polarity to its synonyms and connected words, analyzing changes in polarity for words across different times and domains. In Figure 6, we reproduce the illustration of this algorithm.
Figure 6: Adapted from Hamilton2016InducingDS. The SentiProp algorithm propagates the polarity from seed words based on distance of the connected nodes. Words are connected based on co-occurence statistics.
@@ 4.4 Embeddings
Word embeddings have been used for various tasks, including LSC identification and characterization. This is done by comparing word vectors between two moments in time; the distance between them, and relative to neighboring vectors, can indicate if the word’s sense has changed (Hu2019DiachronicSM).
@@ 4.4.1 Dimension
Ryskina2020WhereNW proposed a statistical approach to identify neologisms based on frequency count and semantic sparsity (measured with embeddings). They claim that words and concepts are subject to supply and demand, so emerging concepts (i.e., neologisms or sense broadening) appear to populate sparse semantic regions until the space becomes uniformly covered.
In Giulianelli2020AnalysingLS, the authors use BERT to create contextualized embeddings of words and then cluster them to obtain usage types. They apply entropy difference, Jensen-Shannon divergence (JSD), and average pairwise distance (APD) as proxies to measure word broadening and narrowing, finding that higher APD and JSD were the best metrics for broadening.
Figure 7: Plot obtained from moss2020detecting. The author represent words as gaussian embeddings, and analyze it’s variance and proximity over time. The word gay increased in variance and got closer to ‘homosexual.’
By analyzing the deviation of a Gaussian word embedding, Yksel2021SemanticCD identified a way to measure narrowing and broadening in word senses. In moss2020detecting, the authors also performed per-sense training of these embeddings. In Figure 7, we observe the distribution of the word ‘gay’ becoming wider and closer to ‘homosexual’.
Lautenschlager2024DetectionON uses XL-LEXEME (Cassotti2023XLLEXEMEWP) clustered embeddings to progressively discover new senses in the historical COHA corpus. They frame the task as Unknown Sense Detection: given an initial dictionary, the goal is to identify unrecorded senses. The system compares a word usage to all its dictionary senses, and if the highest similarity score falls below a threshold, the usage is flagged as a potential new sense.
@@ 4.4.2 Relation
Fonteyn2021AdjustingSA investigate the metaphorization of the concept “to death.” They combined cluster analysis and sentiment analysis to contextualize occurrences of this phrasal verb and detect if metaphorization had taken place. Another approach is to create a training corpus and use it to train a BERT model to predict if a word is metaphorical or literal. The work of Maudslay2022MetaphoricalPD follows this approach and proposes a framework to detect words that started being used metaphorically.

### v5

@@ 2.3 Review Criteria
Table 1: Search terms used to discover articles related to change characterization.
@@
“novel sense change”,“innovative meaning”,“new sense identification”,
“neologism identification”
@@ 3 Corpora for Semantic Change Analysis
Corpora for semantic change can be selected based on differences in time (diachronic studies), domain (academic fields (Schlechtweg et al., 2019)), culture (social media (Carpuat et al., 2013; Shoemark et al., 2019), countries), or dialect (Takamura et al., 2017; Uban et al., 2021), provided the corpora share a sufficient vocabulary for comparison.
In diachronic studies, some methods divide a corpus into discrete time windows, striking a balance between the size of the window and the quantity of data it contains. In contrast, others use a continuous representation of time, providing the timestamp of each text directly to the model (Hu et al., 2019; Rosenfeld and Erk, 2018; Wang et al., 2021). These corpora were designed for specific purposes. For example, corpora with discrete time bins enable a comparative perspective on meaning, while continuous time representations allow for an evolutionary view of the changes.
With respect to characterization, the collected papers show an increasing interest in diachronic corpora, with many proposing handcrafted rules for automatically annotating change categories (Tang et al., 2013). Few works rely on human annotation (Schlechtweg et al., 2017). While these approaches have advantages and drawbacks, the data required to train large representation models is often insufficient from a computational perspective.
Most of the works reviewed here follow a specific formula for diachronic studies, one historical corpus ($C_{1}$) and a more modern corpus ($C_{2}$). Where splitting Corpus of Historical American English (COHA) Davies (2012), a diversified corpus with a span from 1810s to the 2000s, in two corpus is the standard choice. Recent state-of-the-art NLP models require training on billions or even trillions of tokens (He et al., 2021), but diachronic corpora are often on the scale of millions of tokens (Schlechtweg et al., 2020). Given that increasing the amount of historical training data is often infeasible, a viable approach is to develop better algorithms for learning from limited data. In Table 2 we present the table of diachronic corpus in the reviewed works.
@@
Corpus Name Description Statistics (Date, Size) CLMET A collection of diachronic English corpora covering various text types like literary works, letters treatise and others. 34 million words, spanning from 1710-1920, divided into three 70-year sub-periods. COHA A balanced diachronic corpus of American English compiled from diverse document genres. Over 400 million words from about 107,000 documents published from the 1810s to the 2000s. People’s Daily Newspaper A Chinese newspaper corpus used for its faithful recording of contemporary language and events. Spans from 1946 to 2004, with a yearly average of 10,886,017 tokens. Google N-Gram Book The largest available historical corpus, compiled from about 4% of all ever-published books. Over 1TB of text data, containing about 0.3 trillion words and spanning from 1600 to 2009. Kubhist2 A sample of a Swedish historical newspaper archive. From 1790-1830 containing about 71 million tokens. DTA A high-quality collection of historical German texts selected for representativeness and genre balance. 1,022 texts from the period 1741-1900. Lampeter Corpus A collection of British English texts, encompassing various domains such as religion, politics, and law. Approximately one million words from 1640-1740. BNC A collection of primarily written British English sources. One hundred million words, from the late 20th century. NYT Archives A collection of newspaper articles from The New York Times archives. 13 million articles (1851-present). Helsinki Corpus Comprises texts spanning Old English, Middle English, and Early Modern English. The analysis focused on the Middle and Modern English. Approximately 1.1 million words. Old (1̃150), Middle (1150-1500), and Modern (1500-1710) Project Gutenberg The bulk of English language literary works available through the project’s website. 4034 separate documents consisting of over 290 million words.
Even with progress in low-resource learning fields like rapid adaptation (Neubig and Hu, 2018), few-shot learning (Liu et al., 2022), or in-context learning (Min et al., 2022), there is still a significant lack of freely available corpora for semantic change studies (Kutuzov et al., 2018).
@@ 4.1.1 Dimension
To detect changes along the dimension axis ($\mathcal{D}$), Sagi et al. (2009) proposed counting co-occurrences of words and phonemes and measuring the cluster density of the cosine distance on a decomposition of the term-frequency matrix. They assumed that Latent Semantic Analysis is appropriate for examining vectors representing the context of each term’s occurrence and estimated cohesiveness by measuring the density of these context vectors. While they explored the methodology for taxonomic change analysis, they state that the method in fact measures word polysemy.
Tang et al. (2013) proposed a framework for semantic change computation based on the assumption that it is an inherently successive and diachronic process. They distinguished between the word level (overall meaning) and the sense level (individual senses), building their approach around these two notions. A key step involves first identifying a word’s senses in a corpus and then analyzing their diachronic distribution.
Emms and Jayapal (2016) proposed a generative conditional model to detect if a word has acquired a new sense (a semantic neologism). They modeled the word’s sense distribution using Gibbs sampling, applying the model to time-stamped text to infer how the probabilities of a word’s senses change over time. The authors obtain all usages for a word in a specific year ($C_{1}$) from Google N-gram Corpus and compare with the usage frequency for the following next year($C_{2}$), obtaining n-grams that change above a threshold.
Bochkarev et al. (2022) used a neural network to predict whether a word is used as a named entity. With this approach, they constructed a temporal view of word usage by tracking occurrences as nouns versus non-nouns over time, aiming to detect if the word acquired a new meaning.
@@ 4.1.2 Relation
The work of Tang et al. (2013) transforms the problem of relation identification into detecting change patterns in time series data. They analyze semantic change at the word and individual-sense levels, using rules to infer whether a word is used metonymically or metaphorically. A different approach computes the entropy of target words to identify if a word was used in a non-standard context (Schlechtweg et al., 2017). The authors split the usage of words from the Deutsches Textarchiv (DTA) in two time periods (varying by word) and measure the increase in entropy as a proxy for metaphorization. According to the authors, a word’s higher entropy makes it more likely to be used in a newly created metaphorical sense. However, the authors do not differentiate this from the emergence of a new, unrelated sense. They use a corpus of German documents to validate their approach.
@@ 4.1.3 Orientation
Frequency metrics that detect co-occurrences with sets of “good” or “bad” sentiment words have been widely used to infer the amelioration and pejoration of a word over time, especially for neologisms (Cook and Stevenson, 2010; Jatowt and Duh, 2014). Similarly, Buechel et al. (2016) used seed words in German and English to track emotions captured in corpora. They represent these emotions in the Valence-Arousal-Dominance framework rather than simple positive/negative connotation. A limitation of the latter study is that it doesn’t track the evolution of orientation in words.
@@ 4.2.1 Dimension
These approaches assume that changes in the number of topics assigned to a word can indicate changes in dimension. Lau et al. (2012) introduced the idea of novelty as the proportion of a word’s usages corresponding to senses in a focus corpus versus a reference corpus. The goal is to use word sense induction based on topic modeling to detect if a word is used with a new meaning.
Cook et al. (2013) extended this work by quantifying a topic’s usage, estimating that a previously ”culturally silent” topic that becomes frequent in a focus corpus is highly likely to contain neologisms. They introduced the concept of topic relevance to compute a ranked list of potential neologisms.
Figure 4: Figure adapted from Inoue et al. (2022). The stacked bar plots represent the topics obtained over time for the words ‘coach’, ‘record’ and ‘power’ respectively. We can observe new senses emerging and becoming dominant.
Lau et al. (2014) improved the definition of novelty by adding the word’s frequency in the corpus, then proposed combining novelty and relevance to rank neologisms. Frermann and Lapata (2016) introduced SCAN, a Bayesian topic model where topics are preserved across adjacent time periods. Inoue et al. (2022) introduced InfiniteSCAN, an improvement over SCAN. While most topic models rely on a predefined number of topics, InfiniteSCAN uses a logistic stick-breaking process to adapt the number of topics dynamically. This modification allows it to track a varying number of senses, enabling to count each meaning to measure broadening and narrowing, as shown in Figure 4. The authors split Clean COHA in intervals of 20 years to represent the senses.
@@ 4.2.2 Relation
To the best of our knowledge, no methodology for tackling conceptual change with topic-based representation has been published, although a work uses it in a synchronic detection setting (Heintz et al., 2013).
@@ 4.2.3 Orientation
To the best of our knowledge, no methodology approaching orientation change with topic-based representation has been published. Wankhade et al. (2022) use topic models in their workflow to refine data for a sentiment classifier, but topics are not directly used to detect change in feelings.
@@ 4.3.1 Dimension
Comparing the size and structure of a graph over time can provide rich information about a word’s semantic change. Biemann (2006) was among the first to use this approach to detect sense broadening. He created a graph of associated nouns from a corpus and then applied the Chinese whispers algorithm to find clusters. By comparing these clusters over time, he could infer whether a word had acquired a new meaning.
Mitra et al. (2015) analyzed tweets, counting word co-occurrences in bi-grams to build a graph. They ran a graph-clustering algorithm (Chinese Whispers) to determine a set of senses, where each cluster represents a different sense. By comparing these sets of senses over time, they could infer if a word’s senses split or became broader.
Figure 5: Adaptation from Ehmüller et al. (2020). Ego-network, built from word co-occurrence graph, for ‘mouse’. We observe that in 1830 it was used with the sense of ‘weak’ and ‘rat,’ where in 1960 the sense of ‘computer device’ emerged.
Mathew et al. (2017) improved the filtering and sense selection methods of previous works (Mitra et al., 2015; McCarthy et al., 2004; Lau et al., 2014) to better capture sense evolution in Google Books and newspaper datasets. Tahmasebi and Risse (2017) considered a graph of two-word noun phrases as a cluster of meaning, then merged sub-graphs based on Lin similarity (Lin1998AnID). They tracked change in dimension by analyzing how these sub-graphs change over time.
Ehmüller et al. (2020) built a co-occurrence graph, filtering it with point-wise mutual information to remove non-relevant edges, and then created ego-networks to run graph-clustering algorithms. This approach allowed them to build a sense tree—a tree of word usage evolution where new senses are new branches, as shown in Figure 5. Similar to previous work, the authors restricted their analysis to nouns only.
@@ 4.3.2 Relation
Graphs are a good way to represent word relations, and some approaches have used them to detect metonymy usage, e.g., Teraoka (2016); Pedinotti and Lenci (2020). Their models use WordNet or a graph-based distributional model to determine a contextualized representation of word semantics. However, to the best of our knowledge, no work has proposed using graphs at a representation level to compare which words have changed their relations between corpora.
@@ 4.3.3 Orientation
The orientation of words is generally inferred from similarity or proximity to seed words. Many word graphs are constructed for this purpose, and some works (Reagan et al., 2017; Barnes et al., 2021) use these graphs to infer orientation. Hamilton et al. (2016a) used the SentiProp algorithm to propagate a word’s polarity to its synonyms and connected words, analyzing changes in polarity for words across different times and domains. In Figure 6, we reproduce the illustration of this algorithm.
Figure 6: Adapted from Hamilton et al. (2016a). The SentiProp algorithm propagates the polarity from seed words based on distance of the connected nodes. Words are connected based on co-occurence statistics.
@@ 4.4 Embeddings
Word embeddings have been used for various tasks, including LSC identification and characterization. This is done by comparing word vectors between two moments in time; the distance between them, and relative to neighboring vectors, can indicate if the word’s sense has changed (Hu et al., 2019).
@@ 4.4.1 Dimension
Ryskina et al. (2020) proposed a statistical approach to identify neologisms based on frequency count and semantic sparsity (measured with embeddings). They claim that words and concepts are subject to supply and demand, so emerging concepts (i.e., neologisms or sense broadening) appear to populate sparse semantic regions until the space becomes uniformly covered.
In Giulianelli et al. (2020), the authors use BERT to create contextualized embeddings of words and then cluster them to obtain usage types. They apply entropy difference, Jensen-Shannon divergence (JSD), and average pairwise distance (APD) as proxies to measure word broadening and narrowing, finding that higher APD and JSD were the best metrics for broadening.
Figure 7: Plot obtained from Moss (2020). The author represent words as gaussian embeddings, and analyze it’s variance and proximity over time. The word gay increased in variance and got closer to ‘homosexual.’
By analyzing the deviation of a Gaussian word embedding, Yüksel et al. (2021) identified a way to measure narrowing and broadening in word senses. In Moss (2020), the authors also performed per-sense training of these embeddings. In Figure 7, we observe the distribution of the word ‘gay’ becoming wider and closer to ‘homosexual’.
Lautenschlager et al. (2024) uses XL-LEXEME (Cassotti et al., 2023) clustered embeddings to progressively discover new senses in the historical COHA corpus. They frame the task as Unknown Sense Detection: given an initial dictionary, the goal is to identify unrecorded senses. The system compares a word usage to all its dictionary senses, and if the highest similarity score falls below a threshold, the usage is flagged as a potential new sense.
@@ 4.4.2 Relation
Fonteyn and Manjavacas (2021) investigate the metaphorization of the concept “to death.” They combined cluster analysis and sentiment analysis to contextualize occurrences of this phrasal verb and detect if metaphorization had taken place. Another approach is to create a training corpus and use it to train a BERT model to predict if a word is metaphorical or literal. The work of Maudslay and Teufel (2022) follows this approach and proposes a framework to detect words that started being used metaphorically.